Obeyed can slide on a smooth circular wire frame of radius R which is fixed in a vertical plane the beat is displaced slightly from the highest point of the wire flip the speed of the beat subsequently as a function of the angle theta made by the beach with a vertical line is

Grade:11

1 Answers

Khimraj
5 years ago
let angle made with verticle line is \Theta
then
verticle displacement = R – R*cos\Theta = R(1 – cos\Theta)
then applying work energy theorm
mgh = ½*m*v2
v = \sqrt{2gh} = \sqrt{2*g*R(1-cos\Theta )}
Hope it clears. If you like answer then please approve it.